RektbyFaith:

1. One of the tables below contains values that were generated by a linear function. Determine which table, and then write the equation of the linear function represented by the table.

RektbyFaith:

I believe it would be A but I don't know the equation for it

1 attachment
Vocaloid:

not A. remember, for a linear function, x and y must increase/decrease by a constant amount each time

RektbyFaith:

oh I forgot about that

Vocaloid:

for table A: y = 1, 3, 7, 13, 21, 31, 43 it goes +2, then +4, then +6 so this isn't linear

RektbyFaith:

So it would be C

Vocaloid:

good

RektbyFaith:

X= an increase by 2 constantly Y= increase by 5 constantly

Vocaloid:

good using this same pattern, can you try to determine what the y value would be when x = 0?

RektbyFaith:

perhaps -3 ?

RektbyFaith:

nvm,

Vocaloid:

almost

RektbyFaith:

one sec

RektbyFaith:

-4

Vocaloid:

good so if x = 0, y = -4, -4 is your intercept your slope is rise/run = change in y / change in x = (5/2) so your entire equation is y = (5/2)x - 4

RektbyFaith:

Wow, Thank you Vocaloid.

Vocaloid:

unfortunately I have other obligations this afternoon but hopefully somebody can help you if you have other questions ;_;

RektbyFaith:

yeah, I will later, but thank you!